News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 1 December 1999 6.00am EST

During the last 48 hours Chandra observed the GTO target PSR B0656+15 with HRC-S/LETG, X0614+091 with HRC-S/LETG and the HRC-S/LETG calibration source, Cal 83. The observatory passed through the radiation belts and began a post-belt ACIS CTI measurement.

Assessment of the results from the ACIS FEP-0 engineering software patch test performed Nov 26-27 show that no parity errors occurred during the test period. The test would have checked the patch functionality if the events had occurred. Work will now proceed to complete the final version of the patch and uplink to the spacecraft once regression testing is complete.

Observations will continue for the next 48 hours with an HRC-S calibration observation of Cal 83, of NGC 526A, RX J0019.8+2156 with HRC-S, a series of observations of M31 with HRC-S, 3C 58 with HRC-I, NGC 4038 and NGC 4039 with ACIS-S, and of V1416+4446 with ACIS-I.

The spacecraft subsystems continued to operate nominally.





Return to Status Reports